OpenStudy (anonymous):

HELP! PLEASE! THIS IS MATTER OF LIFE AND DEATH!! The rental car charges $25 plus a fixed charge per mile. For the use of 210 miles the rental car was charged $67. What would be the cost of the rental car for 510 miles of use? Is the answer

OpenStudy (anonymous):

lil dramatic..but anyways y=mx+25 (point (210, 67) 67= m(210)+25

OpenStudy (anonymous):

solve for m

OpenStudy (anonymous):

How..?

OpenStudy (anonymous):

Take away 25 from both sides?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

42 = m(210)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

What do you do after that..? divide each side by 210?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

.2 = m

OpenStudy (anonymous):

that's what i got. So now you have a formula to calculate the missing value y=.2x +25

OpenStudy (anonymous):

x represents miles y=price

OpenStudy (anonymous):

y = .2(500) + 25

OpenStudy (anonymous):

its 510 miles (but close)

OpenStudy (anonymous):

y = 100 + 25 y = 125

OpenStudy (anonymous):

its not 500*.2 it is 510*.2, so it would be 102+25 (but you had the right idea)

OpenStudy (anonymous):

No..but apparently, the answer was $two hundred something..said my algebra teacher..

OpenStudy (anonymous):

really i find that hard to believe. If they were able to go 210 miles for 67%, going a little more than double that should not be over 200

OpenStudy (anonymous):

*$67 But anyways. I guess you're right. I hope we got it right..?

OpenStudy (anonymous):

hope so =o

OpenStudy (anonymous):

I thought the equation was y = 25x + 67 why was I wrong?

OpenStudy (anonymous):

from the way the question is written it says that the rental car is charged 25 and a fixed amount per mile. the 25 dollars is not dependent upon the distance traveled